OpenStudy (anonymous):

PLEASE HELP ME!!!! I WILL GIVE A MEDAL AND BECOME A FAN......... IM EXTREMELY CONFUSED AND I HAVE BEEN WORKING ON THIS PROBLEM FOR THE PAST DAY!Interest Problem: Julian deposits money into two different savings accounts. He deposits $600 more into Account B than Account A. After a certain amount of time, Account A has earned $180 at a rate of 4%, and Account B has earned $375 at a rate of 5%. How much did Julian initially deposit into each account? A) What equation is used to solve this problem? What does each variable represent? B) Identify the given information. What do you know about P? C) Enter the given information into the equation. D) Solve the equation. E) State the solution.

OpenStudy (anonymous):

@surjithayer can u help me?

OpenStudy (anonymous):

let the principal in account A=x then principal in account B=x+600 Let time=t \[interest A1=\frac{ x*4*t }{100}=180 ...(1)\] \[Interest A2=\frac{ \left( x+600 \right)*5*t }{100 }=375 ...(2)\] divide (2) by (1) \[\frac{ \left( x+600 \right)*5*t }{100 }*\frac{ 100 }{ 4xt }=\frac{ 375 }{ 100 }\] \[\frac{ 5x+3000 }{4x }=\frac{ 25 }{12 }\] solve and get the value of x

OpenStudy (anonymous):

I still don't understand